kellief59:

melania has two jobs. During the day she works as an office clerk, and in the evening she works as a cashier. Her office job pays her $13 per hour. Her cashier job pays her $9.25 per hour. In one week, Melania worked 38 hours. She earned a total of $434. let c represent the number of hours that Melania works as an office clerk. which equation can be used to find how many hours melania worked in both jobs.

dude:

Uh what are the equation options?

darkknight:

lol, without an equation we can't help you

Mercury:

choices: 13c + 9.25(c – 38) = 434 13c + 9.25(38 – c) = 434 13c + 9.25(434 – c) = 38 9.25c + 13(38 – c) = 434 from the problem: c = # of hours worked as an office clerk she works a total of 38 hours so 38 - c = # of hours worked as a cashier (hourly wage as clerk)(hours worked as clerk) + (hourly wage as cashier)(hours worked as cashier) = total earnings plug in the appropriate #'s and variables to build the equation